# GLGENN: A Novel Parameter-Light Equivariant Neural Networks Architecture Based on Clifford Geometric Algebras

### v1

@@ 1 Introduction
Figure 1: GLGENN is an architecture of neural networks equivariant with respect to any pseudo-orthogonal transformation. Inputs and outputs are represented as multivectors (elements of geometric algebras), which encode various geometric quantities such as scalars, vectors, oriented areas (bivectors) and volumes (trivectors), and higher-dimensional objects (4-vectors, etc.). GLGENN are parameter-light, since they operate in a unified manner across $4$ fundamental subspaces of geometric algebras defined by the grade involution ($\widehat{\;\;}$) and reversion ($\widetilde{\;\;}$); they processes geometric objects in groups with a step size of $4$.
@@
The idea of leveraging GAs to construct pseudo-orthogonal group equivariant neural networks was first proposed by the Amsterdam Machine Learning Lab (University of Amsterdam). Their pioneering work Ruhe et al., 2023 introduced Clifford Group Equivariant Neural Networks (CGENN) and served as the foundation for many subsequent research works. Brehmer et al., 2023 propose Geometric Algebra Transformer (GATr), which incorporates GA into ordinary transformer architecture and outperforms traditional non-geometric baselines in $N$-body modeling and robotic planning. Zhdanov et al., 2024 introduce Clifford-Steerable Convolutional Neural Networks based on CGENN approach, demonstrating superior performance in tasks related to fluid dynamics and relativistic electrodynamics forecasting. Liu et al., 2024 introduce Clifford Group Equivariant Simplicial Message Passing Networks, a method for steerable $\mathrm{E}(n)$-equivariant message passing on simplicial complexes. Pepe et al., 2024 use CGENN to predict protein coordinates to estimate the 3D structure of a protein.
@@ 2.1 Geometric (Clifford) Algebras
The identity element of $C\kern-1.00006pt\ell_{p,q,r}$ is denoted by $e\equiv 1$, the generators are denoted by $e_{a}$, $a=1,\ldots,n$. The generators satisfy the following conditions: $e_{a}e_{b}+e_{b}e_{a}=2\eta_{ab}e$ for any $a,b=1,\ldots,n$, where $\eta=(\eta_{ab})$ is the diagonal matrix with $p$ times $+1$, $q$ times $-1$, and $r$ times 00 on the diagonal in the real case $C\kern-1.00006pt\ell(\mathbb{R}^{p,q,r})$ and $p+q$ times $+1$ and $r$ times 00 on the diagonal in the complex case $C\kern-1.00006pt\ell(\mathbb{C}^{p+q,0,r})$.
Let us consider the subspaces $C\kern-1.00006pt\ell^{k}_{p,q,r}$ of fixed grades $k=0,\ldots,n$. Their elements are linear combinations of the basis elements $e_{a_{1}\ldots a_{k}}:=e_{a_{1}}\cdots e_{a_{k}}$, $a_{1}<\cdots<a_{k}$. The grade-00 subspace can be denoted by $C\kern-1.00006pt\ell^{0}$ without the lower indices $p,q,r$, since it does not depend on the GA’s signature. We have $C\kern-1.00006pt\ell^{k}_{p,q,r}=\{0\}$ for $k<0$ and $k>n$. We can represent any element (multivector) $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum of $n+1$ elements $U=\langle U\rangle_{0}+\cdots+\langle U\rangle_{n},\;\langle U\rangle_{k}\in C\kern-1.00006pt\ell^{k}_{p,q,r}$, $k=0,\ldots,n$. We call any operation of the form
@@
We can represent any element $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum $U=\langle U\rangle_{(0)}+\langle U\rangle_{(1)},\quad\langle U\rangle_{(l)}\in C\kern-1.00006pt\ell^{(l)}_{p,q,r}$, $l=0,1$. We use angle brackets $\langle\cdot\rangle_{(l)}$ to denote the operation of projection of multivectors and sets onto $C\kern-1.00006pt\ell^{(l)}_{p,q,r}$. The grade involution and reversion define four subspaces $C\kern-1.00006pt\ell^{\overline{0}}_{p,q,r}$, $C\kern-1.00006pt\ell^{\overline{1}}_{p,q,r}$, $C\kern-1.00006pt\ell^{\overline{2}}_{p,q,r}$, and $C\kern-1.00006pt\ell^{\overline{3}}_{p,q,r}$ (they are called the subspaces of quaternion types $0,1,2$, and $3$ respectively in Shirokov, 2012a, b):
@@
In other words, $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}:=C\kern-1.00006pt\ell^{k}_{p,q,r}\oplus C\kern-1.00006pt\ell^{k+4}_{p,q,r}\oplus C\kern-1.00006pt\ell^{k+8}_{p,q,r}\oplus\cdots$ for $k=0,1,2,3$. A discussion on the significance of the grade involution, reversion, and these subspaces is provided in Appendix C. Note that the GA $C\kern-1.00006pt\ell_{p,q,r}$ can be represented as a direct sum of the subspaces $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, and viewed as $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded algebra with respect to the commutator and anticommutator (Shirokov, 2018). We can represent any element $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum of $4$ elements: $U=\langle U\rangle_{\overline{0}}+\langle U\rangle_{\overline{1}}+\langle U\rangle_{\overline{2}}+\langle U\rangle_{\overline{3}}$, where $\langle U\rangle_{\overline{m}}\in C\kern-1.00006pt\ell^{\overline{m}}_{p,q,r}$, $m=0,1,2,3$. The action of the grade involution, reversion, and Clifford conjugation on a multivector $U$ of the form above is summarized in Table 1. Note that $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}=C\kern-1.00006pt\ell^{k}_{p,q,r}$, $k=0,1,2,3$, in the cases $n\leq 3$.
@@
$C\kern-0.90005pt\ell^{\overline{k}}_{p,q,r}$ $k=0$ $k=1$ $k=2$ $k=3$
$\widehat{\;\;\;}$ $+$ -- $+$ --
$\widetilde{\;\;\;}$ $+$ $+$ -- --
$\widehat{\widetilde{\;\;\;}}$ $+$ -- -- $+$
@@ 2.2 Lipschitz Groups and Twisted Adjoint Representations
for any $U\in C\kern-1.00006pt\ell_{p,q,r}$ and $T\in C\kern-1.00006pt\ell^{\times}_{p,q,r}$. The representation $\tilde{{\rm ad}}$ has the following properties, which we prove in Lemma E.2 and apply in construction of equivariant layers in Section 3.3. The properties of ${\rm ad}$ and $\check{{\rm ad}}$, are considered in Lemma E.2 as well.
@@ 3.2 Pseudo-orthogonal Groups, Complex Orthogonal Groups, and Generalized Lipschitz Groups Equivariance
[…]
@@ 3.3 Generalized and Ordinary Lipschitz Groups Equivariant Mappings
=$\!\!\!=\begin{cases}\langle T\langle x\rangle_{\overline{m}}T^{-1}\rangle_{\overline{m}}=T\langle x\rangle_{\overline{m}}T^{-1},&\mbox{$m$ is even},\\ \langle\widehat{T}\langle x\rangle_{\overline{m}}T^{-1}\rangle_{\overline{m}}=\widehat{T}\langle x\rangle_{\overline{m}}T^{-1},&\mbox{$m$ is odd},\end{cases}$ (42)
@@
Since $\tilde{\Gamma}^{\overline{1}}_{p,q,r}\subseteq(C\kern-1.00006pt\ell^{(0)\times}_{p,q,r}\cup C\kern-1.00006pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ (Theorem 3.3), $\tilde{{\rm ad}}_{T}$ is multiplicative for any $T\in\tilde{\Gamma}^{\overline{1}}_{p,q,r}$ by Lemma 2.1. The mappings $\psi(x)$ and $\chi(x)$ are $\tilde{\Gamma}^{\overline{1}}_{p,q,r}$-equivariant by Lemma G.4, since they are products of the $\tilde{\Gamma}^{\overline{1}}_{p,q,r}$-equivariant mappings: $x\mapsto\widetilde{x}$, $x\mapsto\widehat{\widetilde{x}}$, and $x\mapsto x$ (Theorem 3.7). ∎
@@ 4.1 Geometric Algebra Data Embeddings
GLGENN, along with other GA-based equivariant neural networks such as CGENN (Ruhe et al., 2023), can be applied to any data representable as multivectors. For example, consider the following task in $\mathbb{R}^{0,3}$. We have data objects $x_{1},\ldots,x_{l}$, where each $x_{i}$ contains a scalar (e.g. $a\in\mathbb{R}$, representing mass, temperature, or a one-hot-encoded feature, etc.), a vector (e.g. $(b,c,d)\in\mathbb{R}^{0,3}$, where $b,c,d\in\mathbb{R}$, representing position, velocity, etc.), and a pseudoscalar (e.g., a signed volume value $f\in\mathbb{R}$). These components can be embedded in GA as $ae+be_{1}+ce_{2}+de_{3}+fe_{123}\in C\kern-1.00006pt\ell_{0,3}$ and processed by our layers. Additional grades beyond 00, $1$, and $3$ can be utilized as well. [^2] Note that in the case of $C\kern-1.00006pt\ell_{0,3}$, the equivariant model proposed in this paper has some connection with equivariant models based on irreducible representations (irreps) of ${\rm O}(3)$ (Han et al., 2022). Generally speaking, these are two different approaches: irreps of ${\rm O}(3)$ follow the tensor approach, whereas the GAs operate with multivectors. However, there are connections between the two approaches: scalars, vectors, bivectors, and trivectors in GA are equivalent to $l=0$ irreps with even parity, $l=1$ irreps with odd parity, $l=1$ irreps with even parity, and $l=0$ irreps with odd parity, respectively. The dimensions are the same ($1$, $3$, $3$, and $1$, respectively), the behavior under rotations/reflections of the coordinate system matches, and the coupling operations (such as scalar multiplication, vector scalar and cross products in various cases) are also the same. Further study of this connection deserves a separate study and is not the topic of the current work, but seems important for building bridges between the two communities. The output can be obtained by projecting the resulting multivectors onto the appropriate subspace based on the task. E.g., projecting onto $C\kern-1.00006pt\ell^{0}$ allows for scalar predictions (which will be both equivariant and invariant), or onto $C\kern-1.00006pt\ell^{1}_{p,q,r}$ if vector predictions are required.
@@ 4.2 Conjugation Operations Layers
where $\phi_{c_{in}k}\in\{-1,1\}$ are optimizable coefficients. The conjugation operation layer is $\tilde{\Gamma}^{1}_{p,q,r}$-equivariant by Theorem 3.8 and contains $l(n+1)$ optimizable parameters. We suggest the following method to make the optimization of conjugation layers with discrete parameters possible. We first apply a linear transformation of the projections of the input multivector as in (44) but with any parameters. Then we round them to $-1$ or $1$ either by directly applying $\phi_{c_{in}k}\mapsto\mbox{sgn}(\phi_{c_{in}k})$ or by firstly applying the sigmoid function, then scaling the value to $[-1,1]$, and applying sign function to the result. Conjugation operations, including $\widehat{\;\;}$, $\widetilde{\;\;}$, and $\widehat{\widetilde{\;\;}}$, play a significant role in GAs and their applications (see Appendix C for a detailed discussion), making these layers highly promising for experiments.
@@ 5 Experiments
Figure 2: Left: ${\rm O}(5,0)$-Regression. Middle: ${\rm O}(5,0)$-Convex Hull, 16 points. The shaded regions depict 95% confidence intervals taken over 5 runs. Right: ${\rm O}(5,0)$-$N$-Body. The shaded regions depict 95% confidence intervals taken over 3 runs.
@@
Table 3: MSE ($\downarrow$) on the ${\rm O}(5,0)$-Convex Hull Experiment ($K$ points). The average convex hull volume in the training dataset for $K=16$, $256$, $512$ points is $\approx 11.4$, $\approx 430.3$, and $\approx 694$, respectively. The number of trainable parameters is as follows: for $K=16$, GLGENN $24.1$K vs. CGENN $58.8$K; for $K=256$, GLGENN $791$K vs. CGENN $1.72$M; for $K=512$, GLGENN $922$K vs. CGENN $1.75$M.
@@
Gap $-1.77$ $-1.53$ 0.1 0.35 0.39 0.52 0.53 $-2268.54$ $-466.53$ $-6188.74$ $-2340.05$
In these experiments, the task is to estimate the volume of a convex hull generated by $K$ points in $\mathbb{R}^{5,0}$ and $\mathbb{R}^{7,0}$ respectively. We first consider the setup with $K=16$ points, as in Ruhe et al., 2023; Liu et al., 2024. The average convex hull volume in a training dataset is $\approx 11.4$. The results, which are presented in Figure 2 (middle) and Table 3, demonstrate that GLGENN either outperform or match CGENN, which itself outperforms ordinary MLP, Geometric Vector Perceptrons (Jing et al., 2021), and Vector Neurons (Deng et al., 2021). In Table 3, MSE for $K=16$ are averaged over 5 runs, number of iterations is the same for both algorithms. Notably, as highlighted Ruhe et al., 2023, CGENN tends to overfit when trained on small datasets. In contrast, GLGENN demonstrate a reduced tendency to overfitting. The reason might be that GLGENN achieve these superior results with more than twice the parameter efficiency, using significantly fewer parameters than CGENN. Note that when CGENN is scaled down to approximately the same size as GLGENN (around $25$K trainable parameters), its performance deteriorates compared to its original configuration (see Table 9). In the case $n=7$, we get similar results (see Appendix I).
@@
$\mathrm{O}(5,0)$-$N$-Body Experiment

### v2

@@ 1 Introduction
Figure 1: GLGENN is an architecture of neural networks equivariant with respect to any pseudo-orthogonal transformation. Inputs and outputs are represented as multivectors (elements of geometric algebras), which encode various geometric quantities such as scalars, vectors, oriented areas (bivectors) and volumes (trivectors), and higher-dimensional objects (4-vectors, etc.). GLGENN are parameter-light, since they operate in a unified manner across 44 fundamental subspaces of geometric algebras defined by the grade involution ($\widehat{\;\;}$) and reversion ($\widetilde{\;\;}$); they processes geometric objects in groups with a step size of 44.
@@
The idea of leveraging GAs to construct pseudo-orthogonal group equivariant neural networks was first proposed by the Amsterdam Machine Learning Lab (University of Amsterdam). Their pioneering work Ruhe et al., 2023 introduced Clifford Group Equivariant Neural Networks (CGENN) and served as the foundation for many subsequent research works. Brehmer et al., 2023 propose Geometric Algebra Transformer (GATr), which incorporates GA into ordinary transformer architecture and outperforms traditional non-geometric baselines in NN-body modeling and robotic planning. Zhdanov et al., 2024 introduce Clifford-Steerable Convolutional Neural Networks based on CGENN approach, demonstrating superior performance in tasks related to fluid dynamics and relativistic electrodynamics forecasting. Liu et al., 2024 introduce Clifford Group Equivariant Simplicial Message Passing Networks, a method for steerable $\mathrm{E}(n)$-equivariant message passing on simplicial complexes. Pepe et al., 2024 use CGENN to predict protein coordinates to estimate the 3D structure of a protein.
@@ 2.1 Geometric (Clifford) Algebras
The identity element of $C\kern-1.00006pt\ell_{p,q,r}$ is denoted by $e\equiv 1$, the generators are denoted by $e_{a}$, $a=1,\ldots,n$. The generators satisfy the following conditions: $e_{a}e_{b}+e_{b}e_{a}=2\eta_{ab}e$ for any $a,b=1,\ldots,n$, where $\eta=(\eta_{ab})$ is the diagonal matrix with $p$ times +1+1, $q$ times 1-1, and $r$ times 0 on the diagonal in the real case $C\kern-1.00006pt\ell(\mathbb{R}^{p,q,r})$ and $p+q$ times +1+1 and $r$ times 0 on the diagonal in the complex case $C\kern-1.00006pt\ell(\mathbb{C}^{p+q,0,r})$.
Let us consider the subspaces $C\kern-1.00006pt\ell^{k}_{p,q,r}$ of fixed grades $k=0,\ldots,n$. Their elements are linear combinations of the basis elements $e_{a_{1}\ldots a_{k}}:=e_{a_{1}}\cdots e_{a_{k}}$, $a_{1}<\cdots<a_{k}$. The grade-0 subspace can be denoted by $C\kern-1.00006pt\ell^{0}$ without the lower indices $p,q,r$, since it does not depend on the GA’s signature. We have $C\kern-1.00006pt\ell^{k}_{p,q,r}=\{0\}$ for $k<0$ and $k>n$. We can represent any element (multivector) $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum of $n+1$ elements $U=\langle U\rangle_{0}+\cdots+\langle U\rangle_{n},\;\langle U\rangle_{k}\in C\kern-1.00006pt\ell^{k}_{p,q,r}$, $k=0,\ldots,n$. We call any operation of the form
@@
We can represent any element $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum $U=\langle U\rangle_{(0)}+\langle U\rangle_{(1)},\quad\langle U\rangle_{(l)}\in C\kern-1.00006pt\ell^{(l)}_{p,q,r}$, $l=0,1$. We use angle brackets $\langle\cdot\rangle_{(l)}$ to denote the operation of projection of multivectors and sets onto $C\kern-1.00006pt\ell^{(l)}_{p,q,r}$. The grade involution and reversion define four subspaces $C\kern-1.00006pt\ell^{\overline{0}}_{p,q,r}$, $C\kern-1.00006pt\ell^{\overline{1}}_{p,q,r}$, $C\kern-1.00006pt\ell^{\overline{2}}_{p,q,r}$, and $C\kern-1.00006pt\ell^{\overline{3}}_{p,q,r}$ (they are called the subspaces of quaternion types $0,1,2$, and 33 respectively in Shirokov, 2012a, b):
@@
In other words, $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}:=C\kern-1.00006pt\ell^{k}_{p,q,r}\oplus C\kern-1.00006pt\ell^{k+4}_{p,q,r}\oplus C\kern-1.00006pt\ell^{k+8}_{p,q,r}\oplus\cdots$ for $k=0,1,2,3$. A discussion on the significance of the grade involution, reversion, and these subspaces is provided in Appendix C. Note that the GA $C\kern-1.00006pt\ell_{p,q,r}$ can be represented as a direct sum of the subspaces $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}$, $k=0,1,2,3$, and viewed as $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$-graded algebra with respect to the commutator and anticommutator (Shirokov, 2018). We can represent any element $U\in C\kern-1.00006pt\ell_{p,q,r}$ as a sum of 44 elements: $U=\langle U\rangle_{\overline{0}}+\langle U\rangle_{\overline{1}}+\langle U\rangle_{\overline{2}}+\langle U\rangle_{\overline{3}}$, where $\langle U\rangle_{\overline{m}}\in C\kern-1.00006pt\ell^{\overline{m}}_{p,q,r}$, $m=0,1,2,3$. The action of the grade involution, reversion, and Clifford conjugation on a multivector $U$ of the form above is summarized in Table 1. Note that $C\kern-1.00006pt\ell^{\overline{k}}_{p,q,r}=C\kern-1.00006pt\ell^{k}_{p,q,r}$, $k=0,1,2,3$, in the cases $n\leq 3$.
@@
$C\kern-0.92505pt\ell^{\overline{k}}_{p,q,r}$ $k=0$ $k=1$ $k=2$ $k=3$
$\widehat{\;\;\;}$ ++ - ++ -
$\widetilde{\;\;\;}$ ++ ++ - -
$\widehat{\widetilde{\;\;\;}}$ ++ - - ++
@@ 2.2 Lipschitz Groups and Twisted Adjoint Representations
for any $U\in C\kern-1.00006pt\ell_{p,q,r}$ and $T\in C\kern-1.00006pt\ell^{\times}_{p,q,r}$. The representation $\tilde{{\rm ad}}$ has the following properties, which we prove in Lemma E.2 and apply in construction of equivariant layers in Section $3.$ The properties of $\rm ad$} and $\check{{\rm ad}}$, are considered in Lemma E.2 as well.
@@ 3.2 Pseudo-orthogonal Groups, Complex Orthogonal Groups, and Generalized Lipschitz Groups Equivariance
$\!\!\!{\rm O}(V,\mathfrak{q})$ $:=\{\Phi:V\rightarrow V:\quad\mbox{$\Phi$ is linear, invertible},$
@@ 3.3 Generalized and Ordinary Lipschitz Groups Equivariant Mappings
$\!\!\!=\begin{cases}\langle T\langle x\rangle_{\overline{m}}T^{-1}\rangle_{\overline{m}}=T\langle x\rangle_{\overline{m}}T^{-1},&\mbox{$m$ is even},\\ \langle\widehat{T}\langle x\rangle_{\overline{m}}T^{-1}\rangle_{\overline{m}}=\widehat{T}\langle x\rangle_{\overline{m}}T^{-1},&\mbox{$m$ is odd},\end{cases}$ (42)
@@
Since $\tilde{\Gamma}^{\overline{1}}_{p,q,r}\subseteq(C\kern-1.00006pt\ell^{(0)\times}_{p,q,r}\cup C\kern-1.00006pt\ell^{(1)\times}_{p,q,r})\Lambda^{\times}_{r}$ (Theorem 3.3), $\tilde{{\rm ad}}_{T}$ is multiplicative for any $T\in\tilde{\Gamma}^{\overline{1}}_{p,q,r}$ by Lemma 2.1. The mappings $\psi(x)$ and $\chi(x)$ are $\tilde{\Gamma}^{\overline{1}}_{p,q,r}$-equivariant by Lemma G.4, since they are products of the $\tilde{\Gamma}^{\overline{1}}_{p,q,r}$-equivariant mappings: $\mapsto\widetilde{x}, xx$$\mapsto\widehat{\widetilde{x}}$, and $x\mapsto x$ (Theorem 3.7). ∎
@@ 4.1 Geometric Algebra Data Embeddings
GLGENN, along with other GA-based equivariant neural networks such as CGENN (Ruhe et al., 2023), can be applied to any data representable as multivectors. For example, consider the following task in $\mathbb{R}^{0,3}$. We have data objects $x_{1},\ldots,x_{l}$, where each $x_{i}$ contains a scalar (e.g. $a\in\mathbb{R}$, representing mass, temperature, or a one-hot-encoded feature, etc.), a vector (e.g. $(b,c,d)\in\mathbb{R}^{0,3}$, where $b,c,d\in\mathbb{R}$, representing position, velocity, etc.), and a pseudoscalar (e.g., a signed volume value $f\in\mathbb{R}$). These components can be embedded in GA as $ae+be_{1}+ce_{2}+de_{3}+fe_{123}\in C\kern-1.00006pt\ell_{0,3}$ and processed by our layers. Additional grades beyond 0, 11, and 33 can be utilized as well. [^2] Note that in the case of $C\kern-1.00006pt\ell_{0,3}$, the equivariant model proposed in this paper has some connection with equivariant models based on irreducible representations (irreps) of ${\rm O}(3)$ (Han et al., 2022). Generally speaking, these are two different approaches: irreps of ${\rm O}(3)$ follow the tensor approach, whereas the GAs operate with multivectors. However, there are connections between the two approaches: scalars, vectors, bivectors, and trivectors in GA are equivalent to $l=0$ irreps with even parity, $l=1$ irreps with odd parity, $l=1$ irreps with even parity, and $l=0$ irreps with odd parity, respectively. The dimensions are the same (11, 33, 33, and 11, respectively), the behavior under rotations/reflections of the coordinate system matches, and the coupling operations (such as scalar multiplication, vector scalar and cross products in various cases) are also the same. Further study of this connection deserves a separate study and is not the topic of the current work, but seems important for building bridges between the two communities. The output can be obtained by projecting the resulting multivectors onto the appropriate subspace based on the task. E.g., projecting onto $C\kern-1.00006pt\ell^{0}$ allows for scalar predictions (which will be both equivariant and invariant), or onto $C\kern-1.00006pt\ell^{1}_{p,q,r}$ if vector predictions are required.
@@ 4.2 Conjugation Operations Layers
where $\phi_{c_{in}k}\in\{-1,1\}$ are optimizable coefficients. The conjugation operation layer is $\tilde{\Gamma}^{1}_{p,q,r}$-equivariant by Theorem 3.8 and contains $l(n+1)$ optimizable parameters. We suggest the following method to make the optimization of conjugation layers with discrete parameters possible. We first apply a linear transformation of the projections of the input multivector as in (44) but with any parameters. Then we round them to 1-1 or 11 either by directly applying $\phi_{c_{in}k}\mapsto\mbox{sgn}(\phi_{c_{in}k})$ or by firstly applying the sigmoid function, then scaling the value to [1,1][-1,1], and applying sign function to the result. Conjugation operations, including $\widehat{\;\;}$, $\widetilde{\;\;}$, and $\widehat{\widetilde{\;\;}}$, play a significant role in GAs and their applications (see Appendix C for a detailed discussion), making these layers highly promising for experiments.
@@ 5 Experiments
Figure 2: Left: ${\rm O}(5,0)$-Regression. Middle: ${\rm O}(5,0)$-Convex Hull, 16 points. The shaded regions depict 95% confidence intervals taken over 5 runs. Right: ${\rm O}(5,0)$-NN-Body. The shaded regions depict 95% confidence intervals taken over 3 runs.
@@
Table 3: MSE ($\downarrow$) on the ${\rm O}(5,0)$-Convex Hull Experiment ($K$ points). The average convex hull volume in the training dataset for $K=16$, $256$, $512$ points is $\approx 11.4$, $\approx 430.3$, and $\approx 694$, respectively. The number of trainable parameters is as follows: for $K=16$, GLGENN 24.124.1K vs. CGENN 58.858.8K; for $K=256$, GLGENN 791791K vs. CGENN 1.721.72M; for $K=512$, GLGENN 922922K vs. CGENN 1.751.75M.
@@
Gap 1.77-1.77 1.53-1.53 0.1 0.35 0.39 0.52 0.53 2268.54-2268.54 466.53-466.53 6188.74-6188.74 2340.05-2340.05
In these experiments, the task is to estimate the volume of a convex hull generated by $K$ points in $\mathbb{R}^{5,0}$ and $\mathbb{R}^{7,0}$ respectively. We first consider the setup with $K=16$ points, as in Ruhe et al., 2023; Liu et al., 2024. The average convex hull volume in a training dataset is $\approx 11.4$. The results, which are presented in Figure 2 (middle) and Table 3, demonstrate that GLGENN either outperform or match CGENN, which itself outperforms ordinary MLP, Geometric Vector Perceptrons (Jing et al., 2021), and Vector Neurons (Deng et al., 2021). In Table 3, MSE for $K=16$ are averaged over 5 runs, number of iterations is the same for both algorithms. Notably, as highlighted Ruhe et al., 2023, CGENN tends to overfit when trained on small datasets. In contrast, GLGENN demonstrate a reduced tendency to overfitting. The reason might be that GLGENN achieve these superior results with more than twice the parameter efficiency, using significantly fewer parameters than CGENN. Note that when CGENN is scaled down to approximately the same size as GLGENN (around 2525K trainable parameters), its performance deteriorates compared to its original configuration (see Table 9). In the case $n=7$, we get similar results (see Appendix I).
@@
$\mathrm{O}(5,0)$-NN-Body Experiment